OpenStudy (anonymous):

simplify 3 ^4√3 -8 ^3√3 +5 ^4√3 +4 ^3√3

OpenStudy (anonymous):

\[3\sqrt[4]{3}-8\sqrt[3]{3}+5\sqrt[4]{3}+4\sqrt[3]{3}\]

OpenStudy (anonymous):

i got -1.121313415....

OpenStudy (anonymous):

all the answers have at least one √

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

its simplifying not solving

OpenStudy (anonymous):

and i simplified it

OpenStudy (anonymous):

ok...can you show me?...

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

give me a minute

OpenStudy (anonymous):

are you sure its a small 3 not a big one?

OpenStudy (anonymous):

for which one?

OpenStudy (anonymous):